# The Dial-a-Ride Problem with Limited Pickups per Trip

### v2

@@
Abstract: The Dial-a-Ride Problem (DARP) is an optimization problem that involves determining optimal routes and schedules for several vehicles to pick up and deliver items at minimum cost. Motivated by real-world carpooling and crowdshipping scenarios, we introduce an additional constraint imposing a maximum number on the number of pickups per trip. This results in the Dial-a-Ride Problem with Limited Pickups per Trip (DARP-LPT). We apply a fragment-based method for DARP-LPT, where a fragment is a partial path. Specifically, we extend two formulations from Rist & Forbes (2021): the Fragment Flow Formulation (FFF) and the Fragment Assignment Formulation (FAF). We establish FFF’s superiority over FAF, both from a theoretical as well as from a computational perspective. Furthermore, our results show that FFF and FAF significantly outperform traditional arc-based formulations in terms of solution quality and time. Additionally, compared to the two existing fragment sets, one with longer partial paths and another with shorter ones, our newly generated fragment sets perform better in terms of solution quality and time when fed into FFF.
Keywords: routing, dial-a-ride problem, limited number of pickups per trip, fragments
@@ 1 Background
DARP has found extensive applications in the sharing mobility (Mourad et al., 2019; Tafreshian et al., 2020) and food delivery sectors. Within the sharing mobility domain, the ride-sharing routing problem represents a variant of DARP. This problem focuses on optimizing routes for drivers, allowing them to efficiently reach their destinations while accommodating passenger pickups and drop-offs, all within the minimum possible time or cost (Dessouky & Mahtab, 2022). Similarly, in the context of online food delivery, the central system needs to make decisions on how to serve submitted requests by efficiently picking up and delivering orders with the goal of minimizing delivery time, which can be seen as an online variant of DARP (Guo et al., 2022).
@@
A frequently disregarded yet vital aspect in this type of routing problems is the limit on the number of pickups per trip. Many studies tend to focus on constraints such as time windows and maximum ride duration, always assuming these can substitute for pickup limits. Nevertheless, explicitly incorporating restrictions on the number of pickups can be crucial for optimizing delivery operations, especially within the shared mobility sector. For instance, in the context of carpooling, each driver typically accommodates a maximum of four passengers per trip, considering vehicle seat capacity and passenger convenience. Service providers such as UberPool [^1] https://www.uberguide.net/how-many-people-can-ride-in-an-uber/, DiDi [^2] https://web.didiglobal.com/au/help-center/how-many-people-are-allowed-on-a-didi-share-ride/, and BlaBlaCar [^3] https://support.blablacar.com/hc/en-gb/articles/360014470360 enforce such restrictions, allowing drivers to pick up no more than two to four passengers on a single trip. Indeed, optimizing a route by initially picking up one passenger and dynamically delivering other passengers, ultimately reaching the first passenger’s destination, is technically feasible. However, private mobility companies often avoid this strategy due to considerations related to passenger comfort. Another influential factor in this consideration is driver preference. In food delivery platforms like Meituan [^4] https://finance.sina.com.cn/tech/2022-05-10/doc-imcwipii9054944.shtml, drivers possess the autonomy to determine the highest number of orders they undertake at one time. Similarly, in the crowdshipping domain, considering drivers’ preferences is crucial as they have the right to set their preferred maximum number for parcel pickups (Arslan et al., 2019). In particular, a company should not anticipate an occasional driver, who might only intend to deliver a few parcels on a sporadic basis or is simply looking to earn some extra income, to handle a high volume of parcels like a full-time staff driver. A similar assumption is found in the concept of limited delivery stops as discussed in Simoni et al. (2020) and Fessler et al. (2022). In this paper, a “customer” can entail more than one person and/or item, but gives rise to only one pickup and two stops (one for pickup and one for delivery).
@@
Since MDARP-LPT and DARP-LPT share the same problem structure, we collectively refer to them as (M)DARP-LPT. To efficiently address (M)DARP-LPT, we propose a unified method with two innovative components: fragment-based formulations and novel fragment generation procedures. Firstly, we tailor the fragment-based formulation of DARP proposed by Rist & Forbes (2021) to accommodate (M)DARP-LPT by modifying the decision variables from the fragment-based format to the fragment-vehicle-based format. This change is also necessary for DARP-LPT, where all homogeneous vehicles start and end at the same depot. Secondly, considering the limited number of pickups, we introduce two novel fragment generation procedures. These procedures create improved fragment sets, enhancing computational efficiency. Traditionally, there are two fragment sets: one with longer partial paths and the other with shorter ones. Our two proposed sets diverge: the first refines the fragment set with longer partial paths through path enumeration, while the second blends longer and shorter partial paths from both traditional fragment sets.
From a practical perspective, this research addresses a commonly overlooked aspect in the dial-a-ride problem: the limited number of pickups per trip. Our developed method offers more tailored and efficient solutions for the problem, thereby contributing to the improvement of carpooling and crowdshipping services. From an academic perspective, this study compares and evaluates various fragment sets for the fragment-based formulation in routing problems, providing essential insights for tackling similar pickup and delivery challenges. Furthermore, our innovative fragment generation method accelerates computations through enumeration; this logic can offer methodological insights for addressing similar routing problems.
The remainder of this article is organized as follows: Section 2 provides a comprehensive literature review. Following that, Section 3 formally defines (M)DARP-LPT, presenting an arc-based formulation and two fragment-based formulations. Section 4 introduces traditional and two novel methods for fragment generation. Then, in Section 5, we analyze the results of our computational experiments. Finally, we draw conclusions in Section 6.
@@ 2 Literature review
Since (M)DARP-LPT is a variant of DARP, we investigate existing research methodologies on related optimization problems, such as the Pickup and Delivery Problem (PDP) and the Dial-a-Ride Problem. PDP involves transporting a set of items from their pickup locations to their respective delivery locations (where multiple items from a single pickup node may be destined for different delivery nodes), with the objective of finding optimal routes while minimizing transportation costs (Savelsbergh & Sol, 1995). DARP is a special case of PDP where each location corresponds to a single item, and every item is associated with a unique pickup and delivery location pair.
@@
Traditional methods, like branch and cut (Cordeau, 2006) and branch and price (Ropke & Cordeau, 2009; Baldacci et al., 2011), have been utilized for PDP and DARP, but recent studies suggest that fragment-based methods exhibit better performance. Here, a fragment is a partial path where only the start and end node have an empty load.
A notable example of the fragment-based approach is the work of Alyasiry et al. (2019), who apply a fragment-based strategy to address the Pickup and Delivery Problem with Time Windows, following the Last-In-First-Out loading strategy. They employ the symbol sequence ``$ppppdddd"$ to denote a full fragment, with “$p$” representing a pickup node and “$d$” indicating a delivery node (this representation is also adopted below). This notation implies that, within the fragment, the last pickup must precede the first delivery. Alyasiry et al. (2019) enumerate all possible paths by combining fragments, thus leading to a fragment-based formulation on the time-space network.
The fragments’ applicability extends to various related problems, retaining the characteristic of empty-loaded start and end nodes but not limited to the specific “$ppppdddd$” structure. Su et al. (2023) effectively compute the minimum excess user ride time in an electric autonomous DARP using the fragment-based method. Additionally, Zhang et al. (2023) tackle a vehicle-customer coordination problem that shares similarities with DARP but incorporates the flexibility for passengers to move in order to meet their drivers faster. To tackle this problem, they design an algorithm based on the fragment-based time-space network. Their procedure not only achieves effective results, but also demonstrates its value for online DARP.
@@
Building on Alyasiry et al. (2019), Rist & Forbes (2021) propose the use of restricted fragments to address the dial-a-ride problem. Since the full fragments in this problem may have the form “$ppdpdpdd$” (as an illustration that the last pickup does not precede the first delivery, this does not imply a universal pattern), their number would be quite high. Nevertheless, Rist & Forbes find that a full fragment with the form “$ppdpdpdd$” can be decomposed into the restricted fragments “$ppd$” +“$pd$”+ “$pdd$”, which start at a pickup node whose predecessor is a delivery node (or depot) and ends at a delivery node whose successor is a pickup node (or depot). Obtaining these restricted fragments involves initially deriving them from fragments conforming to the form “$ppppdddd$” (where the last pickup must precede the first delivery) and then applying a dominance rule to minimize their count. By incorporating temporal information into these restricted fragments, they enable a formulation based on a physical network instead of the time-space network.
While recent literature has made significant contributions to fragment-based methods, other methods may be further improved for (M)DARP-LPT. In particular, alternative methods for fragment generation may result in a higher-quality set of fragments, further enhancing computational performance. Moreover, to the best of our knowledge, no study has conducted a comparison of various fragment types or explored the possibility of combining different fragment categories.
@@ 3 Problem statement and linear formulations
Both MDARP-LPT and DARP-LPT are extensions of DARP, and their major difference from DARP lies in the limitations on the number of pickups per trip. In this section, we begin by providing a formal problem description of DARP-LPT and a NP-hardness proof in Section 3.1. Subsequently, we present several linear formulations for DARP-LPT: an arc-based formulation in Section 3.2 and two fragment-based formulations in Section 3.3. The arc-based formulation closely resembles the one outlined in Cordeau (2006), except for the constraints on the limited number of pickups. The two fragment-based formulations expand on the framework of Rist & Forbes (2021), transitioning from a formulation using fragment-based decision variables to one employing fragment-vehicle-based variables. This adaptation is tailored to our specific requirements for limited number of pickups. Finally, in Section 3.4, we introduce MDARP-LPT, illustrating the differences in its definition and formulations as compared to DARP-LPT.
@@ 3.1 Problem description
DARP-LPT is NP-hard. This can be proven by a reduction from the Capacitated Vehicle Routing Problem (CVRP). CVRP involves arranging a route plan using a fleet of vehicles to visit a set of customers at minimum cost. Each vehicle has a capacity denoted by $C$ (Cornuejols & Harche, 1993; Achuthan et al., 2003). To perform the reduction, in a DARP-LPT instance, we set the maximum number of pickups per trip $L$ equal to the vehicle capacity $C$ and merge each customer’s pickup and delivery nodes into a single node. This creates a problem instance equivalent to CVRP, where each merged node represents a customer with unit size and each vehicle has the capacity $C$ equal to $L$. Since CVRP is known to be NP-hard when $C\geq 3$ (Chen, 2023), this reduction demonstrates that DARP-LPT is also NP-hard. MDARP-LPT can also be proven to be NP-hard using a reduction similar to DARP-LPT.
@@ 3.2 Arc-based formulation of DARP-LPT
The arc-based formulation (ABF), as presented below, is the same as that of Cordeau (2006), except for the constraints (13).
@@
$\sum_{j\in N}f_{ijv}$ $=\sum_{j\in N}f_{(i+n)jv}$ $i\in P,v\in V$ (3)
@@
$\sum_{i\in N}f_{i(2n+1)v}$ $=1$ $v\in V$ (6)
@@
$t_{(n+i)v}-t_{iv}$ $\leq R_{i}$ $i\in P,v\in V$ (11)
$t_{(2n+1)v}-t_{0v}$ $\leq R_{0}$ $v\in V$ (12)
@@
In order to reduce part of the symmetry that is inherent in the formulation, we define $V_{f}=\{1,2,\ldots,\lceil n/L\rceil\}$ as the fixed set of vehicles that must be used. This definition arises from the observation that, if a vehicle can deliver at most $L$ customers, a minimum of $\lceil n/L\rceil$ vehicles are needed. Therefore, we add $f_{0(2n+1)v}=0$ for each fixed vehicle $v\in V_{f}$.

### v3

@@
Abstract
The Dial-a-Ride Problem (DARP) is an optimization problem that involves determining optimal routes and schedules for several vehicles to pick up and deliver items at minimum cost. Motivated by real-world carpooling and crowdshipping scenarios, we introduce an additional constraint imposing a maximum number on the number of pickups per trip. This results in the Dial-a-Ride Problem with Limited Pickups per Trip (DARP-LPT). We apply a fragment-based method for DARP-LPT, where a fragment is a partial path. Specifically, we extend two formulations from Rist & Forbes [18]: the Fragment Flow Formulation (FFF) and the Pickup-Space Fragment Formulation (PSFF). Furthermore, our results show that PSFF outperforms FFF, which in turn surpasses traditional arc-based formulations in both solution quality and computational efficiency. Additionally, we compare several existing fragment sets that differ in the length of their partial paths and find that the sets with shorter partial paths yield the best solution times when used with PSFF. In addition, we propose a new mixed fragment set, which is useful when the sets with longer partial paths become too large. In such cases, it yields the lowest CPU time.
keywords:
routing, dial-a-ride problem, limited number of pickups per trip, fragments
journal: European Journal of Operational Research
@@ 1 Background
DARP has found extensive applications in the sharing mobility [15, 25] and food delivery sectors. Within the sharing mobility domain, the ride-sharing routing problem represents a variant of DARP. This problem focuses on optimizing routes for drivers, allowing them to efficiently reach their destinations while accommodating passenger pickups and drop-offs, all within the minimum possible time or cost [9]. Similarly, in the context of online food delivery, the central system needs to make decisions on how to serve submitted requests by efficiently picking up and delivering orders with the goal of minimizing delivery time, which can be seen as an online variant of DARP [14].
@@
A frequently disregarded yet vital aspect in this type of routing problems is the limit on the number of pickups per trip. Many studies tend to focus on constraints such as time windows and maximum ride duration, always assuming these can substitute for pickup limits. Nevertheless, explicitly incorporating restrictions on the number of pickups can be crucial for optimizing delivery operations, especially within the shared mobility sector. For instance, in the context of carpooling, each driver typically accommodates a maximum of four passengers per trip, considering vehicle seat capacity and passenger convenience. Service providers such as UberPool [^1] https://www.uberguide.net/how-many-people-can-ride-in-an-uber/, DiDi [^2] https://web.didiglobal.com/au/help-center/how-many-people-are-allowed-on-a-didi-share-ride/, and BlaBlaCar [^3] https://support.blablacar.com/hc/en-gb/articles/360014470360 enforce such restrictions, allowing drivers to pick up no more than two to four passengers on a single trip. Indeed, optimizing a route by initially picking up one passenger and dynamically delivering other passengers, ultimately reaching the first passenger’s destination, is technically feasible. However, private mobility companies often avoid this strategy due to considerations related to passenger comfort. Another influential factor in this consideration is driver preference. In food delivery platforms like Meituan [^4] https://finance.sina.com.cn/tech/2022-05-10/doc-imcwipii9054944.shtml, drivers possess the autonomy to determine the highest number of orders they undertake at one time. Similarly, in the crowdshipping domain, considering drivers’ preferences is crucial as they have the right to set their preferred maximum number for parcel pickups [3]. In particular, a company should not anticipate an occasional driver, who might only intend to deliver a few parcels on a sporadic basis or is simply looking to earn some extra income, to handle a high volume of parcels like a full-time staff driver. A similar assumption is found in the concept of limited delivery stops as discussed in Simoni et al. [22] and Fessler et al. [10]. In this paper, a “customer” can entail more than one person and/or item, but gives rise to only one pickup and two stops (one for pickup and one for delivery).
@@
Since MDARP-LPT and DARP-LPT share the same problem structure, we collectively refer to them as (M)DARP-LPT. To efficiently address (M)DARP-LPT, we propose a unified approach featuring two distinct fragment-based formulations and evaluate different types of fragment sets as input. Both formulations are adapted from the DARP framework of Rist & Forbes [18]. The first converts fragment-based decision variables into a fragment–vehicle–based form, a necessary adjustment even for (single-depot) DARP-LPT. The second introduces a novel pickup-space fragment formulation, extending the network with a pickup count dimension, similar to time in the time-space formulation, and which uses pickup-space decision variables. As for fragment sets: traditional methodologies employ several distinct sets, each with fragments (partial paths) of different levels of length. We also introduce a novel mixed fragment set, combining fragments of different lengths to reduce the total number while preserving useful ones as much as possible.
From a practical perspective, this research addresses a commonly overlooked aspect in the dial-a-ride problem: the limited number of pickups per trip. Our developed method offers more tailored and efficient solutions for the problem, thereby contributing to the improvement of carpooling and crowdshipping services. From an academic perspective, this study proposes novel extended fragment-based formulations and conducts a comprehensive comparison and evaluation of multiple fragment sets for routing problems. Our findings offer valuable insights for addressing various problems that incorporate pickup and delivery. Additionally, this study is the first to explore the combination of different fragment types for constructing the fragment set used in the formulations, providing valuable insights for routing problems.
The remainder of this article is organized as follows: Section 2 provides a comprehensive literature review. Following that, Section 3 formally defines (M)DARP-LPT, presenting an arc-based formulation, two fragment-based formulations, and a path-based formulation. Section 4 introduces traditional and novel methods for fragment generation. Then, in Section 5, we analyze the results of our computational experiments. Finally, we draw conclusions in Section 6.
@@ 2 Literature review
Since (M)DARP-LPT is a variant of DARP, we investigate existing research methodologies on related optimization problems, such as the Pickup and Delivery Problem (PDP) and the Dial-a-Ride Problem. PDP involves transporting a set of items from their pickup locations to their respective delivery locations (where multiple items from a single pickup node may be destined for different delivery nodes), with the objective of finding optimal routes while minimizing transportation costs [21]. DARP is a special case of PDP where each location corresponds to a single item, and every item is associated with a unique pickup and delivery location pair.
@@
Traditional methods, like branch and cut [6] and branch and price [19, 4], have been utilized for PDP and DARP, but recent studies suggest that fragment-based methods exhibit better performance. Here, a fragment is a partial path where only the start and end node have an empty load.
A notable example of the fragment-based approach is the work of Alyasiry et al. [2], who apply a fragment-based strategy to address the Pickup and Delivery Problem with Time Windows, following the Last-In-First-Out loading strategy. Examples can be represented as $p1p2p3d3d2d1$ or $p1p2d2p3d3d1$, where “$p$” denotes a pickup node, “$d$” denotes a delivery node, and the number indicates the customer index. This representation will also be used throughout the remainder of the paper. By combining fragments, Alyasiry et al. [2] are able to enumerate all feasible paths; based on this, they construct a fragment-based formulation on a time-space network. The fragments’ applicability extends to various related problems, retaining the characteristic of empty-loaded start and end nodes but not limited to the specific Last-In-First-Out setting. Subsequently, these fragments are referred to as “full fragments”. In this context, Su et al. [24] effectively compute the minimum excess user ride time in an electric autonomous DARP using the fragment-based method.
@@
Building on Alyasiry et al. [2], Rist & Forbes [18] propose the use of restricted fragments to address the dial-a-ride problem. Since the full fragments in this problem don’t necessarily adhere to the Last-In-First-Out property, their number would be quite high. However, Rist & Forbes find that a full fragment with, for instance, the form “$ppdpdpdd$” can be decomposed into the restricted fragments “$ppd$” +“$pd$”+ “$pdd$”, which start at a pickup node whose predecessor is a delivery node (or depot) and ends at a delivery node whose successor is a pickup node (or depot). Note that “$ppdpdpdd$” is merely an illustration where the last pickup does not necessarily precede the first delivery, this does not imply a universal pattern. Rather, it is intended to contrast with the form “$ppppdddd$”, where the last pickup must precede the first delivery. Obtaining these restricted fragments involves initially deriving them from fragments conforming to the form “$ppppdddd$” and then applying a dominance rule to minimize their count. Unlike Alyasiry et al. [2], Rist & Forbes [18] formulate the model on a physical network instead of a time-space network. Time-related constraints are handled purely via callbacks, yet the computational burden remains modest.
Subsequently, Rist & Forbes [17] introduced the concept of extended full fragments for the Selective Dial-a-Ride Problem (SDARP), a variant of DARP that aims to maximize passenger service. An extended full fragment is defined as the concatenation of a full fragment followed by an empty arc. Notably, the instances used in SDARP feature stricter time windows and shorter ride times than those in standard DARP. These tighter constraints motivate the use of a time-space network and help ensure a manageable number of extended full fragments. Rist & Forbes’s [2022] computational results demonstrate that extended full fragments outperform existing fragment types for such tightly constrained instances. Similarly, Zhang et al. [26] adopt a similar definition of extended full fragments, where each fragment is formed by concatenating an empty arc followed by a full fragment. They tackle a vehicle-customer coordination problem that shares similarities with DARP but incorporates the flexibility for passengers to move in order to meet their drivers faster. In line with Rist & Forbes [17], the test instances in Zhang et al. [26] feature even tighter time windows and ride time constraints (e.g., 1.2 to 1.5 times the delivery time limit of an ideal detour-free trip), which further contributes to keeping the number of fragments manageable. To tackle this problem, they design an algorithm based on a time-space fragment-based network. Their procedure not only achieves effective results, but also demonstrates its value for online DARP.
While recent literature has made significant contributions to fragment-based methods, none can be directly applied to handle the limited number of pickups in our problem statement. Additionally, different fragment types may result in varying computational performance. To the best of our knowledge, no existing study has systematically compared different fragment types or explored the potential of combining multiple fragment categories.
Another recent development is the event-based formulation for DARP [12, 11], where an event represents a vehicle’s state as defined by its location and onboard passengers. The event-based formulation is quite similar to the traditional arc-based formulation of Cordeau [6], but in the event-based model the graph is constructed with events as nodes and feasible transitions as arcs. This design implicitly enforces capacity, precedence, and pairing constraints, thereby reducing the number of explicit constraints and improving computational efficiency compared to traditional arc-based models. Nevertheless, events inherently provide less routing information than fragments, as noted by Gaul et al. [11]. Given the fairly superior performance of the fragment-based method on the benchmark instances of [18], we only concentrate on fragment-based strategies in this study.
@@ 3 Problem statement and linear formulations
Both MDARP-LPT and DARP-LPT are extensions of DARP, and their major difference from DARP lies in the limitations on the number of pickups per trip. In this section, we begin by providing a formal problem description of DARP-LPT and an NP-hardness proof in Section 3.1. Subsequently, we present several linear formulations for DARP-LPT: an arc-based formulation in Section 3.2, two fragment-based formulations in Section 3.3, and a path-based formulation in Section 3.4. The arc-based formulation closely resembles the one outlined in Cordeau [6], except for the constraints on the limited number of pickups. The two fragment-based formulations build upon the framework of Rist & Forbes [18]. The first formulation extends the framework by transitioning from fragment-based decision variables to fragment–vehicle-based ones. The second extends the network into a pickup–space representation, introducing pickup–space fragment-based variables to account for pickup constraints. This adaptation is tailored to our specific requirements for limited number of pickups. The path-based formulation presented in Section 3.4 adopts a set partitioning structure, as in Ropke & Cordeau [19]. Finally, in Section 3.5, we introduce MDARP-LPT, illustrating the differences in its definition and formulations as compared to DARP-LPT.
@@ 3.1 Problem description
DARP-LPT is NP-hard. This can be proven by a reduction from the Capacitated Vehicle Routing Problem (CVRP). CVRP involves arranging a route plan using a fleet of vehicles to visit a set of customers at minimum cost. Each vehicle has a capacity denoted by $C$ [7, 1]. To perform the reduction, in a DARP-LPT instance, we set the maximum number of pickups per trip $L$ equal to the vehicle capacity $C$ and merge each customer’s pickup and delivery nodes into a single node. This creates a problem instance equivalent to CVRP, where each merged node represents a customer with unit size and each vehicle has the capacity $C$ equal to $L$. Since CVRP is known to be NP-hard when $C\geq 3$ [5], this reduction demonstrates that DARP-LPT is also NP-hard. MDARP-LPT can also be proven to be NP-hard using a reduction similar to DARP-LPT.
@@ 3.2 Arc-based formulation of DARP-LPT
The arc-based formulation (ABF), as presented below, is the same as that of Cordeau [6], except for the constraints (13).
@@
$\sum_{j\in N}f_{ijv}$ $=\sum_{j\in N}f_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i+n,j,v}}$ $i\in P,v\in V$ (3)
@@
$\sum_{i\in N}f_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,2n+1,v}}$ $=1$ $v\in V$ (6)
@@
$t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n+i,v}}-t_{iv}$ $\leq R_{i}$ $i\in P,v\in V$ (11)
$t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2n+1,v}}-t_{0v}$ $\leq R_{0}$ $v\in V$ (12)
@@
In order to reduce part of the symmetry that is inherent in the formulation, we define $V_{f}=\{1,2,\ldots,\lceil n/L\rceil\}$ as the fixed set of vehicles that must be used. This definition arises from the observation that, if a vehicle can deliver at most $L$ customers, a minimum of $\lceil n/L\rceil$ vehicles are needed. Therefore, we add $f_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0,2n+1,v}}=0$ for each fixed vehicle $v\in V_{f}$.